OpenStudy (anonymous):

Which expressions, when evaluated with x = –9, are negative in value? Choose all answers that are correct –2x – 8 –26 – 2x 35 + 3x –3x + (–40) will medal :]

OpenStudy (anonymous):

–2x – 8==== x = –9 then 10 –26 – 2x==== x = –9 then -8 35 + 3x==== x = –9 then 8 –3x + (–40)==== x = –9 then -13

OpenStudy (anonymous):

thanks you